INVESTIGATION OF NUCLEAR SPIN CONVERSION OF METHANE IN SOLID HYDROGEN

\underline{MIZUHO FUSHITANI} and TAKAMASA MOMOSE, Division of Chemistry, Graduate School of Science, Kyoto University, Kyoto 606-8502, Japan.

We have studied nuclear spin conversion of CH_{4} and CD_{4} trapped in solid H_{2} and solid D_{2} by Fourier-transform infrared spectroscopy. It has been shown that the rotational quantum number J of methane is still a good quantum number in solid hydrogen. The effective rotational constant has been found to be about 90 % of the gas phase value. The vibrationrotation absorption of both CH_{4} and CD_{4} exhibited time-dependent intensity changes at liquid He temperatures. These changes are interpreted to be a result of the I = 1 → I = 2 nuclear spin conversion which accompanies the J = 1 → J = 0 rotational relaxation. The conversion rate is on the order of 10^{−3} min^{−1}, but the rate depends on the temperature of the crystal. Mechanisms of the nuclear spin conversion in solid hydrogen will be discussed.